# Recurrent Diffusion for Large-Scale Parameter Generation

### v1

@@ Appendix A Discussion with More Related Works
    Additionally, the parameters generated by our model can directly achieve almost peak performance without any finetuning. And the generation process is entirely synthesized from noise, eliminating the need for a few prompt examples that are trained for a few epochs, as required by the $S_{\text{KDE}}$ [61] sampling method.
@@
A.2 Details and limitations of G.pt.
A primary limitation of G.pt [51] is the training data collection cost. By default, they collect 23 million checkpoints to train the parameter generator. Besides, they only evaluate the effectiveness of G.pt on small architectures, such as a low-dimensional MLP layer or a Convolutional layer with limited channels. The maximum number of generated parameters does not exceed 10,000.
A.3 Details and limitations of p-diff.
P-diff [69] directly flattens all parameters into a single-dimensional vector, disregarding the inter-layer parameter relationships. Furthermore, p-diff faces challenges in scaling up to large-scale parameter generation.
Appendix B Experimental Settings and More Results
@@ B.1 Training recipe
In this section, we provide detailed training recipes and supplementary information. The number of parameters generated by our approach ranges from approximately 3K to 200M. The significant disparity necessitates different training settings. Generally, as the number of parameters increases, the learning process becomes more challenging, requiring higher training costs, particularly for generating parameters beyond 50 million. Therefore, our training settings are divided into two categories: the default setting and the setting for parameters exceeding 50 million, as is shown in Tab. 14.
@@
Diffusion batch size: In our approach, the diffusion model is shared across all tokens. Typically, all tokens can be fed as a single batch into the diffusion model for training. However, in practice, we randomly select a subset of tokens from a long sequence for training, rather than feeding all parts at once. This approach significantly reduces memory usage without compromising performance. The “diffusion batch size” in Tab. 14 refers to the number of tokens fed into the diffusion model during a single training iteration.
training setting params. count $<$ 50M params. count $>$ 50M
@@
Table 14: Training recipe in detail. The amount of parameters that need to be generated is denoted as params. count
B.2 Datasets
@@
Classification
  • ImageNet-1k [12] is a large-scale visual database for visual object recognition research. It contains over 1 million images across 1000 categories and is widely used for training and benchmarking deep learning models.
  • CIFAR-10 [32] dataset consists of 60,000 32$\times$32 colorful images in 10 different classes. It is commonly used for training machine learning and computer vision algorithms, providing a standard benchmark for image classification task.
Semantic segmentation
  • ADE20K [77] is a dataset for semantic segmentation and scene parsing, containing over 20,000 images annotated with pixel-level labels for 150 object categories. It is used to train models to understand and segment various objects and scenes in an image, making it valuable for applications in autonomous driving, robotics, and image editing.
Instance segmentation & Object detection
  • COCO [39] dataset is a large-scale object detection, segmentation, and captioning dataset. It contains over 330,000 images, with more than 200,000 labeled instances across 80 object categories. COCO is widely used for training and evaluating models in object detection, segmentation, and image captioning tasks.
Commonsense reasoning:
  • BoolQ [9]: Yes/no questions based on natural passages.
  • PIQA [3]: Questions about physical tasks and actions.
  • SIQA [56]: Questions about social interactions.
  • HellaSwag [75]: Choosing the correct ending for stories.
  • ARC [10]: Multiple-choice science questions.
  • OBQA [45]: Questions requires multi-step reasoning, commonsense knowledge, and rich text comprehension.
B.3 The detailed structure of recurrent diffusion
@@
Details of recurrent model.
By default, the recurrent model consists of two Mamba layers [19]. As the increasing of parameters to generate, we need a larger recurrent model to capture the information in these parameters. The size of the recurrent model is mainly determined by the token size, which varies according to the number of parameters to be generated. Based on the token size, we categorize our model into four versions: Tiny, Small, Base, and Large.
Details of diffusion model.
Following p-diff [69], our diffusion model adopts a one-dimensional convolutional architecture. The parameters of the diffusion model are significantly fewer than those of the recurrent model. We feed the prototypes from the recurrent model as conditions into the diffusion model by directly adding them to the feature map.
@@
B.4 More results of Section 5
@@
Figure 7: Principal Component Analysis (PCA) visualization of the classification head. The figures demonstrate the presence of an inherent structure in the high-dimensional parameter space and highlight our method’s effectiveness in capturing this structure for unseen tasks.
Figure 8: Examples of RPG-generated models guided by binary embeddings from a large language model (Qwen2.5-3B [73]), demonstrating neural network parameter generation conditioned by natural language. For the first example, we give the LLM a prompt: ”Give me a model to select all living things.” With the binary embedding provided by the LLM, our RPG then generates a ViT-Tiny classifier. After that, We use images in CIFAR-10 to evaluate the model’s accuracy. The model should classify ”bird”, ”cat”, ”deer”, ”dog”, ”frog”, and ”horse” to the positive class, which we used as the ground truth. The result is 97.1%. The second example follows the same process.
PCA visualization of classification head parameters.
@@
To demonstrate the application scenarios of our model, we utilize large language model to generate binary embeddings to guide RPG in generating corresponding classification models. We provide two example results in Fig. 8. This experiment demonstrates our method’s capability to generate neural network parameters based on natural language guidance, highlighting the potential applications of our method.
B.5 Training memory cost analysis
@@
We conduct experiments to show the relationship between GPU memory and generated parameter counts in Fig. 9(b). In previous methods, the relationship between GPU memory consumption and the number of parameters in the generated model was quadratic [58] or directly proportional [69]. This limits their practicality and application range. In contrast, our approach demonstrates remarkable efficiency: with equivalent GPU memory usage, it can generate models with 34 to 960 times more parameters compared to previous methods.
B.6 Inference memory cost and sampling time
(b) Visualization of GPU memory v.s. parameter counts. Our method can generate much more parameters than existing approaches e.g. $\textit{\text{S}}_{\textit{\text{KDE30}}}$ [58] using a single NVIDIA H100 80G GPU.
B.7 Parameter sensitivity v.s. performance
B.8 Why not auto-regression?
best avg. min.
original 75.2 74.9 74.7
auto-regressive fail fail fail
non-auto-regressive 75.0 74.8 74.6

### v2

@@ A.2 Discussion with Hyper-Representations
    Additionally, the parameters generated by our model can directly achieve almost peak performance without any finetuning. And the generation process is entirely synthesized from noise, eliminating the need for a few prompt examples that are trained for a few epochs, as required by the $S_{\text{KDE}}$ (Schürholt et al., 2024) sampling.
@@
A.3 Details and limitations of G.pt
A primary limitation of G.pt (Peebles et al., 2022) is the training data collection cost. By default, they collect 23 million checkpoints to train the parameter generator. Besides, they only evaluate the effectiveness of G.pt on small architectures, such as a low-dimensional MLP layer or a Convolutional layer with limited channels. The maximum number of generated parameters does not exceed 10,000.
A.4 Details and limitations of p-diff
P-diff (Wang et al., 2024) directly flattens all parameters into a 1D vector, disregarding the inter-layer parameter relationships. Furthermore, p-diff faces challenges in scaling up to large-scale parameter generation.
A.5 More related works
Diffusion models.
Diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021) gain increasing popularity in recent years, due to their superiority in image generation. Ever since its advent, many works have been done focusing on improving the generation quality and efficiency of diffusion models. For generation quality, Rombach et al. (2022) propose to conduct diffusion in the latent space, enabling high-resolution image synthesis. Peebles & Xie (2023) leverage the transformer (Vaswani, 2017) to explore scalability of diffusion models, proving the possibility of generating higher quality images by increasing model size. As for efficiency problem, efficient samplers (Song et al., 2020; Lu et al., 2022; Song et al., 2023), efficiency models (Fang et al., 2023; So et al., 2024; Yang et al., 2023), and global acceleration approaches (Ma et al., 2024; Pan et al., 2024) are proposed to increase diffusion models’ efficiency. These methods facilitate generating high quality images with less computational and/or memory cost. Although diffusion models for image generation have achieved great success, improving quality and efficiency in large-scale parameter generation remains to be explored.
Appendix B Experimental Setting and Other Details
@@ B.1 Training recipe
In this section, we provide detailed training recipes and supplementary information. The number of parameters generated by our approach ranges from approximately 3K to 200M. The significant disparity necessitates different training settings. Generally, as the number of parameters increases, the learning process becomes more challenging, requiring higher training costs, particularly for generating parameters beyond 50 million. Therefore, our training settings are divided into two categories: the default setting and the setting for parameters over 50 million, shown in Tab. 12.
@@
Diffusion batch size: In our approach, the diffusion model is shared across all tokens. Typically, all tokens can be fed as a single batch into the diffusion model for training. However, in practice, we randomly select a subset of tokens from a long sequence for training, rather than feeding all parts at once. This approach significantly reduces memory usage without compromising performance. The ‘diffusion batch size’ in Tab. 12 refers to the number of tokens fed into the diffusion model during a single training iteration.
training setting #params. $<$ 50M #params. $>$ 50M
@@
Table 12: Training recipe in detail. The amount of parameters that need to be generated is denoted as #params.
B.2 Description of datasets
@@
Classification: ImageNet-1k (Deng et al., 2009) is a large-scale visual database for visual object recognition research. It contains over 1 million images across 1000 categories and is widely used for training and benchmarking deep learning models. CIFAR-10 (Krizhevsky & Hinton, 2009) dataset consists of 60,000 32$\times$32 colorful images in 10 different classes. It is commonly used for training machine learning and computer vision algorithms, providing a standard benchmark for image classification task.
Semantic segmentation: ADE20K (Zhou et al., 2017) is a dataset for semantic segmentation and scene parsing, containing over 20,000 images annotated with pixel-level labels for 150 object categories. It is used to train models to understand and segment various objects and scenes in an image, making it valuable for applications in autonomous driving, robotics, and image editing.
Instance segmentation & object detection: COCO (Lin et al., 2014) dataset is a large-scale object detection, segmentation, and captioning dataset. It contains over 330,000 images in various resolutions, with more than 200,000 labeled instances across 80 object categories. COCO is widely used for training and evaluating models in object detection, segmentation, and image captioning tasks.
Commonsense reasoning: BoolQ (Clark et al., 2019): Yes/No questions based on natural passages. PIQA (Bisk et al., 2020): Questions about physical tasks and actions. SIQA (Sap et al., 2019): Questions about social interactions. HellaSwag (Zellers et al., 2019): Choosing the correct ending for stories. ARC (Clark et al., 2018): Multiple-choice science questions. OBQA (Mihaylov et al., 2018): Questions requires multi-step reasoning, commonsense knowledge, and rich text comprehension.
B.3 Detailed structure of recurrent diffusion
@@
Details of recurrent model.
By default, the recurrent model consists of two Mamba layers (Gu & Dao, 2023). As the increasing of parameters to generate, we need a larger recurrent model to capture the information in these parameters. The size of the recurrent model is mainly determined by the token size, which varies according to the number of parameters to be generated. Based on the token size, we categorize our model into four versions: Tiny, Small, Base, and Large.
Details of diffusion model.
Following p-diff (Wang et al., 2024), our diffusion model adopts a 1D convolutional architecture. The parameters of the diffusion model are significantly fewer than those of the recurrent model. We feed the prototypes from the recurrent model as conditions into the diffusion model by directly adding them to the feature map.
The setting of our main experiments.
In our main experiments, ViT-Base, ConvNeXt-Large, ADE20K, COCO, and DoRA rank 64 used the setting of RPG-Large. CNN (s) and CNN (m) used the setting of RPG-Tiny. And all other experiments used the setting of RPG-Base.
Appendix C Additional Experimental Results
C.1 Effectiveness of permutation state
The effect of permutation state.
collected runs org. w/o permu. state w/ permu. state
1 88.2 88.0 88.1
3 88.3 fail 88.1
10 88.5 fail 88.2
C.2 More results of Section 4
@@
Efficiency comparison.
epoch from scratch finetune RPG initialization + finetune
0 50.0 52.9 94.4
1 61.4 90.3 96.8
5 69.1 96.3 97.3
10 74.9 97.1 97.5
50 86.7 97.7 97.8
PCA visualization of classification head parameters.
@@
Figure 7: Principal Component Analysis (PCA) visualization of the classification head. The figures demonstrate the presence of an inherent structure in the parameter space and highlight our method’s effectiveness in capturing this structure for unseen tasks.
Figure 8: Illustration of RPG-generated models guided by binary embeddings from a large language model (Qwen2.5-3B (Yang et al., 2024)), demonstrating neural network parameter generation conditioned by natural language.
       prompt       expected embedding       acc. (%)
       Give me a model to select all living things.       0,0,1,1,1,1,1,1,0,0       98.7
       Find all vehicles that operate on roads.       0,1,0,0,0,0,0,0,0,1       95.9
       Classify all mammals.       0,0,0,1,1,1,0,1,0,0       97.6
       Select all man-made objects.       1,1,0,0,0,0,0,0,1,1       97.7
       Find all things that are both man-made and can operate on water.       0,0,0,0,0,0,0,0,1,0       98.4
       Select all animals that can be regarded as pets.       0,0,0,1,0,1,0,0,0,0       96.8
       Select all things that can fly.       1,0,1,0,0,0,0,0,0,0       87.7
       Find all animals with fur.       0,0,1,1,1,1,0,1,0,0       70.4
       Select all pets commonly found in households.       0,0,1,1,0,1,0,0,0,0       83.3
       Identify all cold-blooded animals.       0,0,0,0,0,0,1,0,0,0       98.9
Table 17: The table demonstrates examples of generating from abstract prompts using LLM and RPG. We are able to achieve high accuracy on abstract prompts. The expected embedding is used for evaluating the accuracy.
(b) Visualization of GPU memory v.s. parameter counts. Our method can generate much more parameters than existing approaches e.g. $\textit{\text{S}}_{\textit{\text{KDE30}}}$ (Schürholt et al., 2022a) using a single NVIDIA H100 80G GPU.
@@
To demonstrate the application scenarios of our model, we utilize large language model to generate binary embeddings to guide RPG in generating corresponding classification models. For the first example in Fig 8, we give the LLM a prompt: ‘Give me a model to select all living things.’ With the binary embedding provided by the LLM, our RPG then generates a ViT-Tiny classifier. After that, We use images in CIFAR-10 to evaluate the model’s accuracy. The model should classify ‘bird’, ‘cat’, ‘deer’, ‘dog’, ‘frog’, and ‘horse’ to the positive class, which we used as the ground truth. The result is 97.1%. Some other examples are in Tab 17. This experiment demonstrates our method’s capability to generate neural network parameters based on natural language guidance, highlighting the potential applications of our method.
C.3 Training memory cost analysis
@@
We conduct experiments to show the relationship between GPU memory and generated parameter counts in Fig. 9(b). In previous methods, the relationship between GPU memory consumption and the number of parameters in the generated model was quadratic (Schürholt et al., 2022a) or directly proportional (Wang et al., 2024). This limits their practicality and application range. In contrast, our approach demonstrates remarkable efficiency: with equivalent GPU memory usage, it can generate models with 34 to 960 times more parameters compared to previous methods.
C.4 Inference memory cost & sampling time
@@
C.5 Parameter sensitivity v.s. performance
C.6 Details of trained checkpoint collection
This section mainly supplements the collection of checkpoints in Section 3. First, we obtain the pre-trained models, which either come from model libraries (such as timm (Wightman, 2019)) or are trained by ourselves. Then, we fine-tune the full model for one epoch, and save 50 checkpoints during this epoch uniformly. In Tab 14, the term collected runs refers to the number of times the entire process, from pre-training to fine-tuning and saving, is repeated. This is done without fixing the seed, resulting in checkpoints from entirely different permutations.
C.7 Why not auto-regression?
best avg. min.
original 75.2 74.9 74.7
auto-regressive fail fail fail
non-auto-regressive 75.0 74.8 74.6